# ExplainReduce: Summarising local explanations via proxies

### v1

@@ Appendix C Fidelity of the proxy sets as a function of subsample size
Figure 9 shows the complete fidelity plots as a function of subsample size. Each row corresponds to a different dataset, and each column represents a local model generation method. A general trend emerges, indicating improved fidelity with increasing subsample size.
Figure 9: Fidelity of the reduction algorithms as a function of subsample size. The black dashed line indicates the performance of the full set of local explanations.
Appendix D Coverage of the proxy sets as a function of proxy number
Figure 10 shows the complete results of the coverages of the proxy sets. Each row corresponds to a different dataset, and each column represents a local model generation method. A general trend emerges, indicating broader coverage with increasing $k$.
Figure 10: Coverage of the reduction algorithms as a function of proxy set size $k$.
Appendix E Stability of the proxy sets as a function of proxy number
Figure 11 shows the complete results of the stabilities of the proxy sets. Each row corresponds to a different dataset, and each column represents a local model generation method. A general trend emerges, indicating more stable performance with increasing $k$.
Figure 11: Instability of the reduction algorithms as a function of the proxy set size $k$.
Appendix F Full sensitivity analysis
@@
Many of the reduction methods outlined in this paper require the user to select an error threshold $\varepsilon$, with some additionally requiring a minimum coverage constraint $c$. Since $\varepsilon$ is a parameter in computing coverage, its value can significantly impact the performance of the ExplainReduce procedure. In this section, we study the sensitivity of the procedure to these hyperparameters.
@@
As the error tolerance $\varepsilon$ is defined in comparison to the absolute loss, the exact value of the parameter is task- and data-dependent. In this section, we set the error tolerance as a quantile of the training loss of the underlying closed-box function $f$, i.e., $\varepsilon=q(\bm{L}_{BB},p)$, where $q$ denotes the percentile function, $\bm{L}_{BB,i}=\ell(f(\bm{x}_{i}),\bm{y}_{i})$ and $p$ is the percentile value.
@@
As Fig. 12 shows, the reduction methods that require these parameters produce results with nearly identical fidelity across a wide range of both $c$ and $\varepsilon$. The test fidelity of the proxy set for all tested reduction methods exhibits low variance when the error tolerance is set anywhere between the 10th and 50th percentiles of the closed-box training loss values. Additionally, the loss-minimising reduction method with soft coverage produces similar fidelity results across the entire tested range of minimum coverage values, though it often fails to meet the coverage constraint at higher values. It can be concluded that the reduction methods are not sensitive with respect to these hyperparameters. For simplicity, we suggest using $c=0.8$ and $\varepsilon=q(\bm{L}_{BB},0.2)$ as default arguments. When not stated otherwise, these are also the $c$ and $\varepsilon$ values used in other experiments in this paper.
Figure 12: Sensitivity of the reduction methods as a function of error tolerance and minimum coverage in case of the const min loss algorithm.

### v2

@@ Appendix D Fidelity of the proxy sets as a function of subsample size
Figures 21-32 show the complete fidelity plots as a function of subsample size at proxy set size $k=5$. Each Figure corresponds to a different dataset, and each column represents a local model generation method. A general trend emerges, indicating improved fidelity with increasing subsample size.
Figure 21: Adult dataset reduction function comparison as a function of $n$ on fidelity.
Figure 22: Air Quality dataset reduction function comparison as a function of $n$ on fidelity.
Figure 23: Churn dataset reduction function comparison as a function of $n$ on fidelity.
Figure 24: Gas Turbine dataset reduction function comparison as a function of $n$ on fidelity.
Figure 25: Higgs dataset reduction function comparison as a function of $n$ on fidelity.
Figure 26: Jets dataset reduction function comparison as a function of $n$ on fidelity.
Figure 27: Life Expectancy dataset reduction function comparison as a function of $n$ on fidelity.
Figure 28: QM9 dataset reduction function comparison as a function of $n$ on fidelity.
Figure 29: Spam dataset reduction function comparison as a function of $n$ on fidelity.
Figure 30: Sythetic dataset reduction function comparison as a function of $n$ on fidelity.
Figure 31: Telescope dataset reduction function comparison as a function of $n$ on fidelity.
Figure 32: Vehicles dataset reduction function comparison as a function of $n$ on fidelity.
Appendix E Sensitivity analysis
@@
Many of the reduction methods outlined in this paper require the user to select an error threshold $\varepsilon$, with the Balanced algorithm requiring an additional trade-off parameter $\lambda$. Since $\varepsilon$ is a parameter in computing coverage, its value can significantly impact the performance of the ExplainReduce procedure. In this section, we study the sensitivity of the procedure to these hyperparameters.
@@
As the error tolerance $\varepsilon$ is defined in comparison to the absolute loss, the exact value of the parameter is task- and data-dependent. In this section, we set the error tolerance as a quantile of the training loss of the underlying closed-box function $f$, i.e., $\varepsilon=q(\bm{L}_{BB},p)$, where $q$ denotes the percentile function, $\bm{L}_{BB,i}=\ell(f(\bm{x}_{i}),\bm{\hat{y}}_{i})$ and $p$ is the percentile value.
@@
As Fig. 33 shows, the Balanced algorithm produces results with nearly identical fidelity across a wide range $\varepsilon$ at proxy set size $k=6$ and an initial set of $n=500$ local explanations. The test fidelity of the proxy set for all tested reduction methods exhibits low variance when the error tolerance is set anywhere between the 10th and 50th percentiles of the closed-box training loss values. Similarly, the effect of the trade-off parameter $\lambda$ is negligible at proxy set size $k=6$ (initial explanation set of $n=500$ explanations), as shown in Figures 34 and 35. For simplicity, we suggest using $\lambda=0.5$ and $\varepsilon=q(\bm{L}_{BB},0.3)$ as default arguments. When not stated otherwise, these are also the $\lambda$ and $\varepsilon$ values used in other experiments in this paper.
Figure 33: Test set fidelity as a function of error tolerance for the Balanced algorithm.
Figure 34: Test set fidelity as a function of trade-off parameter $\lambda$ for the Balanced algorithm.
Figure 35: Coverage as a function of trade-off parameter $\lambda$ for the Balanced algorithm.
Appendix F Comparing ExplainReduce to other model aggregation methods
Figures 36-47 show the test and train set fidelities, coverage and stability of ExplainReduce, using the Balanced algorithm, against submodular pick [ribeiro2016], GLocalX [setzu2021glocalx], the IP-based aggregation method by [li2022optimal] and a single global linear explainer. Except for GlocalX, which has a slight edge in the Telescope dataset (Fig. 46), our method consistently exceeds or at least matches the performance of other methods.
Figure 36: Adult dataset global comparison
Figure 37: Air Quality dataset global comparison
Figure 38: Churn dataset global comparison
Figure 39: Gas Turbine dataset global comparison
Figure 40: Higgs dataset global comparison
Figure 41: Jets dataset global comparison
Figure 42: Life Expectancy dataset global comparison
Figure 43: QM9 dataset global comparison
Figure 44: Spam dataset global comparison
Figure 45: Synthetic dataset global comparison
Figure 46: Telescope dataset global comparison
Figure 47: Vehicles dataset global comparison
Appendix G Performance of greedy algorithms
In section 4.1, we discussed the performance of greedy approximation algorithms used to solve Problems 13. As we have seen in previous sections, the performance of the greedy and analytical coverage-optimising algorithms is almost equal as a function of the proxy set size and the initial local explanation set size. As mentioned in the section above, solving for the exact loss-minimising proxy set is a computationally challenging problem which requires $\binom{m}{k}$ comparisons in the worst case. Hence, the analysis in this section was limited to $n=m=100,k=5$ for computational tractability. Table 2 shows how the greedy approximation algorithms perform comparably to the analytically optimal variants. For the max coverage algorithm, experimental approximation ratios are substantially better than the worst-case bound described in Sect. 4.1. Furthermore, as we saw in the previous sections, while on the train set, the greedy heuristics do not reach the optimal solution fidelity-wise. However, on the test, the differences between the exact solution and the greedy heuristics vanish. The Balanced algorithm shows particularly poorer performance compared to the greedy min loss algorithm, but it generalises better on the test set while achieving good coverage performance as well. Overall, we can conclude that the greedy approximations offer a computationally feasible approach to generating proxy sets.
Coverage $\uparrow$
Dataset G. Max $c$ A. ratio Balanced A. ratio Max $c$
Air Quality $0.82\pm 0.18$ $0.97\pm 0.04$ $0.8\pm 0.17$ $0.96\pm 0.04$ $0.84\pm 0.18$
Gas Turbine $0.86\pm 0.17$ $0.99\pm 0.01$ $0.86\pm 0.18$ $0.99\pm 0.03$ $0.87\pm 0.18$
Jets $0.86\pm 0.17$ $0.99\pm 0.02$ $0.86\pm 0.17$ $0.99\pm 0.02$ $0.87\pm 0.18$
QM9 $0.78\pm 0.17$ $0.98\pm 0.03$ $0.78\pm 0.17$ $0.97\pm 0.04$ $0.8\pm 0.18$
Train fidelity $\downarrow$
Dataset G. Min Loss A. ratio Balanced A. ratio Min Loss
Air Quality $0.36\pm 0.35$ $1.25\pm 0.21$ $0.47\pm 0.43$ $1.85\pm 0.91$ $0.32\pm 0.35$
Gas Turbine $0.21\pm 0.23$ $8\pm 0.9 0.22±0$$2\pm 0.29$ $1.75\pm 0.81$ $0.17\pm 0.23$
Jets $0.02\pm 0.02$ $1.4\pm 0.4$ $0.02\pm 0.03$ $1.49\pm 0.37$ $0.01\pm 0.02$
QM9 $0.53\pm 0.46$ $1.25\pm 0.21$ $0.59\pm 0.47$ $1.43\pm 0.43$ $0.48\pm 0.47$
Test fidelity $\downarrow$
Dataset G. Min Loss A. ratio Balanced A. ratio Min Loss
Air Quality $0.4\pm 0.29$ $1.11\pm 0.22$ $0.47\pm 0.37$ $1.28\pm 0.33$ $0.37\pm 0.28$
Gas Turbine $0.31\pm 0.17$ $1.2\pm 0.22$ $0.33\pm 0.24$ $1.17\pm 0.16$ $0.28\pm 0.18$
Jets $0.02\pm 0.01$ $1.09\pm 0.18$ $0.02\pm 0.02$ $1.11\pm 0.17$ $0.02\pm 0.01$
QM9 $1.19\pm 0.13$ $01\pm 0.08 1.1$$14\pm 0.13$ $0.97\pm 0.07$ $1.18\pm 0.11$
Table 2: Comparison of the training fidelity and coverage, alongside the observed approximation ratios (A. ratio), between the analytical and approximate optimisation algorithms when applied to $n=100$ explanations produced with lime and with $k=5$ proxies. G. Max $c$ corresponds to the greedy max coverage algorithm, while G. Min Loss and G. Min Loss (min $c$) correspond to the greedy min loss and Balanced algorithms, respectively.
Appendix H Formalisation of the reduction algorithms
H.1 Max Coverage
Problem 4 (Max $k$-Coverage Variant).
Given a dataset $\mathcal{D}=\{(\bm{x}_{j},\bm{y}_{j})\}_{j=1}^{n}$, a set of candidate models $\bm{G}=\{g_{1},\dots,g_{m}\}$, and a loss threshold $\varepsilon>0$, find a subset $\bm{S}\subseteq\bm{G}$ with $|\bm{S}|\leq k$ that maximizes the number of covered samples.
Definition 1 (Coverage Set).
For each model $g_{i}\in\bm{G}$, we define its coverage set $C_{i}\subseteq\mathcal{D}$ as the set of points for which the loss is within the threshold $\varepsilon$:
$C_{i}=\{(\bm{x}_{j},\bm{y}_{j})\in\mathcal{D}\mid\ell(g_{i}(\bm{x}_{j}),\bm{y}_{j})\leq\varepsilon\}$ (14)
Definition 2 (Objective Function).
The total coverage of a subset of models $\bm{S}\subseteq\bm{G}$ is given by the set function $f:2^{\bm{G}}\to\mathbb{Z}_{\geq 0}$:
$f(\bm{S})=\left|\bigcup_{g_{i}\in\bm{S}}C_{i}\right|$ (15)
Theorem 1.
The greedy algorithm for Problem 4 yields a solution $\bm{S}_{greedy}$ such that:
$f(\bm{S}_{greedy})\geq\left(1-\left(\frac{k-1}{k}\right)^{k}\right)f(\bm{S}^{*})\geq\left(1-\frac{1}{e}\right)f(\bm{S}^{*})$ (16)
where $S^{*}$ is the optimal subset of size $k$.
Proof.
The problem is a direct mapping to the Max $k$-Cover problem. We observe the following properties of the objective function $f(\bm{S})$:
  1. 1.
    Monotonicity: For any $\bm{A}\subseteq\bm{B}\subseteq\bm{G}$, $f(\bm{A})\leq f(\bm{B})$, as the union of additional sets cannot decrease the cardinality of covered elements.
  2. 2.
    Submodularity: For any $\bm{A}\subseteq\bm{B}\subseteq\bm{G}$ and $g\in\bm{G}\setminus B$, the marginal gain satisfies $f(\bm{A}\cup\{g\})-f(\bm{A})\geq f(\bm{B}\cup\{g\})-f(\bm{B})$. This follows from the property of set unions where an element already covered by $\bm{A}$ is necessarily covered by $\bm{B}$.
By the results established in nemhauser1978analysis for the maximization of monotone submodular functions under cardinality constraints, the greedy selection strategy—which iteratively picks $g=\arg\max_{g\in\bm{G}}(f(\bm{S}\cup\{g\})-f(\bm{S}))$—achieves the stated lower bound. ∎
H.2 Min Loss
Problem 5 (Min Loss via Improvement Maximization).
Given a dataset $\mathcal{D}$, find a subset $S\subseteq\bm{G}$ with $|S|=k$ that maximizes the total improvement function:
$f(\bm{S})=\mathcal{L}_{base}-\mathcal{L}(\bm{S})$ (17)
where $\mathcal{L}(\bm{S})=\frac{1}{n}\sum_{j=1}^{n}\min_{g_{i}\in S}\ell(g_{i}(\bm{x}_{j}),\bm{\hat{y}}_{j})$ and $\mathcal{L}_{base}$ is the loss of the baseline configuration.
Lemma 1 (Supermodularity of Loss).
The function $\mathcal{L}:2^{\bm{G}}\to\mathbb{R}_{\geq 0}$ defined in Eq. (17) is supermodular.
Proof.
A function $\mathcal{L}$ is supermodular if for all $\bm{A}\subset\bm{B}\subset\bm{G}$ and $v\notin\bm{B}$:
$\mathcal{L}(\bm{A}\cup\{v\})-\mathcal{L}(\bm{A})\leq\mathcal{L}(\bm{B}\cup\{v\})-\mathcal{L}(\bm{B})$ (18)
Let $L_{j}(\bm{S})=\min_{g_{i}\in S}\ell(g_{i}(\bm{x}_{j}),\bm{y}_{j})$. Since $\mathcal{L}(\bm{S})=\frac{1}{n}\sum_{j}L_{j}(\bm{S})$, and the sum of supermodular functions is supermodular, it suffices to prove the property for a single point $j$. Let $\ell_{i}=\ell(g_{i}(\bm{x}_{j}),\bm{y}_{j})$. We evaluate the marginal change $\Delta(\bm{S},v)=L_{j}(\bm{S}\cup\{v\})-L_{j}(\bm{S})$:
$\Delta(\bm{S},v)=\min(\min_{i\in\bm{S}}\ell_{i},\ell_{v})-\min_{i\in\bm{S}}\ell_{i}=\min(0,\ell_{v}-\min_{i\in\bm{S}}\ell_{i})$ (19)
Since $\bm{A}\subset\bm{B}$, it follows that $\min_{i\in\bm{A}}\ell_{i}\geq\min_{i\in\bm{B}}\ell_{i}$. Let $M_{\bm{A}}=\min_{i\in\bm{A}}\ell_{i}$ and $M_{\bm{B}}=\min_{i\in\bm{B}}\ell_{i}$. Then $M_{\bm{A}}\geq M_{\bm{B}}$. The marginal changes are:
$\Delta(\bm{A},v)$ $=\min(0,\ell_{v}-M_{\bm{A}})$
$\Delta(\bm{B},v)$ $=\min(0,\ell_{v}-M_{\bm{B}})$
Since $M_{\bm{A}}\geq M_{\bm{B}}$, then $M_{\bm{A}}\leq\min(0,\ell_{v}-M_{\bm{B}})$, satisfying (18). ∎
Lemma 2 (Submodularity of Loss Reduction).
The improvement function $f(\bm{S})$ is a non-negative, monotone submodular function.
Proof.
Let $h(\bm{S})=-\mathcal{L}(\bm{S})$. We previously established that $\mathcal{L}(\bm{S})$ is supermodular; therefore, $h(\bm{S})$ is submodular [mccormick2005submodular].
  1. 1.
    Submodularity: $f(\bm{S})=\mathcal{L}_{base}+h(\bm{S})$. Since adding a constant does not change the second-order differences of a set function, $f(\bm{S})$ inherits the submodularity of $h(\bm{S})$.
  2. 2.
    Monotonicity: For any $g\in\bm{G}\setminus\bm{S}$, the marginal gain is $\Delta f(g|\bm{S})=f(\bm{S}\cup\{g\})-f(\bm{S})=\mathcal{L}(\bm{S})-\mathcal{L}(\bm{S}\cup\{g\})$. Since $\min_{i\in S\cup\{g\}}\ell_{i}\leq\min_{i\in S}\ell_{i}$, it follows that $\mathcal{L}(\bm{S}\cup\{g\})\leq\mathcal{L}(\bm{S})$, making the marginal gain $\Delta f(g|\bm{S})\geq 0$.
  3. 3.
    Non-negativity: Provided $\mathcal{L}_{base}\geq\mathcal{L}(\bm{S})$ for all $\bm{S}$ (which holds if $\mathcal{L}_{base}$ is the loss of the empty set or a single-model baseline), $f(\bm{S})\geq 0$.
Theorem 2.
The greedy algorithm (reduce) selecting $g=\arg\max\Delta\mathcal{L}(g|S)$ yields a solution $S_{g}$ such that:
$\mathcal{L}_{base}-\mathcal{L}(\bm{S}_{g})\geq\left(1-\frac{1}{e}\right)(\mathcal{L}_{base}-\mathcal{L}(\bm{S}^{*}))$ (20)
where $S^{*}$ is the optimal subset of size $k$.
Proof.
By Lemma 2, $f(\bm{S})$ satisfies the conditions of nemhauser1978analysis. The greedy algorithm maximizes the marginal improvement at each step, ensuring the (11/e)(1-1/e) approximation ratio relative to the maximum possible improvement. ∎
H.3 Formalization of Joint Utility Maximization
Problem 6 (Balanced Utility Maximization).
Find a subset $\bm{S}\subseteq\bm{G}$ with $|\bm{S}|=k$ that maximizes the joint utility function:
$U(\bm{S})=\lambda\cdot C(\bm{S},\varepsilon)+(1-\lambda)\cdot R(\bm{S})$ (21)
where $C(\bm{S},\varepsilon)$ is the coverage function and $R(\bm{S})=\frac{\mathcal{L}_{base}-\mathcal{L}(\bm{S})}{\mathcal{L}_{base}}$ is the normalized loss reduction.
Lemma 3 (Preservation of Submodularity).
The joint utility function $U(\bm{S})$ is monotone submodular for any $\lambda\in[0,1]$.
Proof.
Consider the two components of $U(\bm{S})$:
  1. 1.
    Coverage $C(\bm{S},\varepsilon)$: As established in Problem 4, this is a standard set-cover function, which is monotone and submodular.
  2. 2.
    Loss Reduction $R(\bm{S})$: Let $\mathcal{L}(\bm{S})$ be the supermodular loss function from Problem 2. Then $\mathcal{L}(\bm{S})$ is submodular. Since $R(\bm{S})=\frac{\mathcal{L}_{base}}{\mathcal{L}_{base}}-\frac{\mathcal{L}(\bm{S})}{\mathcal{L}_{base}}=1-\frac{1}{\mathcal{L}_{base}}\mathcal{L}(\bm{S})$, $R(\bm{S})$ is a linear transformation of a submodular function with a positive scaling of the negative term. Thus, $R(\bm{S})$ is submodular.
Since $\lambda$ and $(1-\lambda)$ are non-negative, and the sum of submodular functions is submodular, $U(\bm{S})$ is submodular. Monotonicity follows because adding a model can only increase coverage and decrease (or maintain) loss. ∎
Theorem 3.
The greedy algorithm $\bm{S}_{g}$ for Problem 3 satisfies:
$U(\bm{S}_{g})\geq\left(1-\frac{1}{e}\right)U(\bm{S}^{*})$ (22)
where $\bm{S}^{*}$ is the optimal subset of size $k$.
Proof.
This is a direct application of the result by nemhauser1978analysis. Since $U(\bm{S})$ is a monotone submodular function with $U(\emptyset)=0$ (assuming $\mathcal{L}(\emptyset)=\mathcal{L}_{base}$), the greedy strategy of picking $g=\arg\max_{g\notin S}(U(\bm{S}\cup\{g\})-U(\bm{S}))$ achieves the (11/e)(1-1/e) lower bound. ∎